# Efficient wireless charging of a quantum battery

### v1

@@ IV.3 The case of three charging units
When three qubits in the reservoir are treated as the charger ($n=3$) and the fourth one as a QB, the behaviors of $\mathcal{E}$, $\mathcal{E}_{i}$, and $\mathcal{E}_{c}$ are also structurally similar to those showed in Figs. 2 and 7, so we do not display them here. The charging time $\bar{t}$ and the ergotropy $\bar{\mathcal{E}}$ charged on the QB at $t=\bar{t}$ also show qualitatively the similar $R$ dependence to those showed in Figs. 4 and 8, so we also do not display them here. We only point out here that compared to those achieved with the two charging units, $\lambda\bar{t}$ for both the scenarios I and II can be further slightly shortened, while $\bar{\mathcal{E}}$ can also be enhanced to some extent.
Figure 10: Comparison of $\bar{\mathcal{E}}$ charged on the QB with different number $n$ of charging units. The left (right) two panels are plotted with the chargers and QB being in the initial states $|\psi\rangle_{c_{1}\ldots c_{n}}$ and $|0\rangle$ ($|1\rangle^{\otimes n}$ and $|\varphi\rangle_{e_{1}}$), respectively. The first five points in each panel are plotted with $R=0.1$, 1, 2, 3 and 4, respectively.
Before ending to this section, we provide some further comparison about the charging of a QB by using different number of charging units. First, we present in Fig. 10 a comparison of the ergotropy, where we have denoted by ¯$\bar{\mathcal{E}}_{n=1}$ the ergotropy charged on a QB with $n=1$, and likewise for ¯$\bar{\mathcal{E}}_{n=2}$ and ¯$\bar{\mathcal{E}}_{n=3}$. For the scenario I, from Fig. 10(a) one can see that if $c_{1,2}=1$, ¯$\bar{\mathcal{E}}_{n=2}>\bar{\mathcal{E}}_{n=1}$ when $R\lesssim 10.34$, and this region expands with a decrease in $c_{1,2}$. This, together with Fig. 10(c), shows that if the charger energy is not maximal initially, one can enhance the ergotropy of the QB by using multiple charging units. As for the scenario II, from Fig. 10(b) one can observe that the region in which ¯$\bar{\mathcal{E}}_{n=2}>\bar{\mathcal{E}}_{n=1}$ shrinks with the increase of $e_{1}$. Although Fig. 10(d) shows that the ergotropy can be slightly enhanced for $n=3$, the enhancement is limited. So in contrast to the scenario I, the performance of a QB can be enhanced by using multiple charging units only in the not very strong coupling regime. We have also compared the ergotropy charged on the initially mixed battery state $\rho_{\mathrm{ba}}=\mathrm{diag}\{e_{1},1-e_{1}\}$ with different number of charging units, and a qualitatively similar phenomenon to that showed in Fig. 10 is observed.
@@
where $\Delta\bar{\mathcal{E}}$ is the net ergotropy charged on the QB, $\Delta E_{\mathrm{ch}}$ is the energy output from the charger, while $\rho_{\mathrm{ch}}(\bar{t})$ denotes the state of the charger at time $\bar{t}$, and likewise for $\rho_{\mathrm{ch}}(0)$. Based on this definition we performed numerical calculations and the result is shown in Fig. 11. Clearly, $P_{\bar{\mathcal{E}}}$ can be noticeably enhanced by increasing the coupling strength. In particular, for the scenario I with a single charging unit, if the QB is in its ground state initially (cf. the circles in the left column of Fig. 11), $P_{\bar{\mathcal{E}}}$ approaches 1 for large enough $R$. Moreover, for the case of not very weak coupling, $P_{\bar{\mathcal{E}}}$ decreases when $n$ increases from 1 to 3, whereas this may not always be the case in the weak coupling regime. For the scenario II, as depicted in Fig. 11(c) and (e), if a single charging unit is considered, $P_{\bar{\mathcal{E}}}$ also approaches 1 when $R$ approaches infinity, but in this case its increase with the increase of $R$ is slower than that for the scenario I. When the two and three charging units are considered, although $P_{\bar{\mathcal{E}}}$ also increases with the increase of $R$, the asymptotic value at $R\rightarrow\infty$ is obviously less than 1. This implies that a considerable amount of energy output from the charger is leaked into the reservoir in this situation.
@@ V Summary and discussion
In summary, we have studied the wireless charging of a QB via $n$ ($n\geqslant 1$) charging units, where both the charger and the QB interact with a structured bosonic reservoir (which may be implemented, e.g., by an electromagnetic field inside a lossy cavity) and there is no direct coupling between them. Inspired by the fact that in realistic situation, the charger energy may be not maximal and one rarely wait until a battery runs out before charging, we considered two scenarios of wireless charging, that is, the scenario I in which the charger is in a superposition of the excited and ground states and the QB is empty (i.e., fully discharged) initially, and the scenario II in which the charger is in its excited state and the QB has residual ergotropy (i.e., partially discharged) initially. For both the scenarios, we considered the ergotropy, which quantifies the maximum amount of work a QB could supply during unitary cycles, as a figure of merit for comparing the charging performance. Our results showed that the charging time $\bar{t}$, defined as the interaction time at which the QB reaches its dynamical maximum, decreases with the increase of the coupling strength. The charging time $\bar{t}$ also slightly shortens with an increase in the number of the charging units. The ergotropy $\bar{\mathcal{E}}$ charged on the QB at $t=\bar{t}$, on the other hand, increases with the increasing coupling strength. Specifically, if a single charging unit is used, then in the strong coupling regime, the ergotropy $\bar{\mathcal{E}}$ for the scenario I increases with the increasing amount of the initial energy in the charger, while that for the scenario II is not very sensitive to the residual ergotropy in the QB. In particular, the QB is almost fully charged when the coupling strength is strong enough. In the weak and moderate coupling regimes, however, the case will be different. The incoherent and coherent contributions to ergotropy are also different in different coupling regimes. Moreover, compared to the case of a single charging unit, the ergotropy $\bar{\mathcal{E}}$ charged by two and three charging units is decreased in the strong coupling regime, whereas in the weak and moderate coupling regimes, it can be enhanced to some extent; in particular, such an enhancement is pronounced for the scenario I in which the charger energy is not maximal, and this implies that in this situation the multiple charging units could compensate to some extent the lack of energy in a single charging unit. Finally, we have also considered the charging efficiency which quantifies the proportion of output energy that can be converted into extractable work in the QB. It was found that it also increases with the increase of the coupling strength. If a single charging unit is considered, this efficiency approaches 1 when $R\rightarrow\infty$, while for that of the two and three charging units, it is significantly decreased.
@@
While we considered in this paper the charging of a single-cell QB via $n$ charging units, one may also be concerned with the general case in which a $m$-cell QB is charged via $n$ charging units. Although it is hard to obtain a general result for this case due to the limited computing resource, we performed numerical calculations with $n+m\leqslant 4$ and the results showed that for $m=2$ and 3, the dynamical behavior of $\mathcal{E}$ is structurally similar to that for a single-cell QB, and the maximum ergotropy $\bar{\mathcal{E}}$ charged on a QB can also be noticeably enhanced by increasing the coupling strength, but now it cannot be fully charged, even if $R$ approaches infinity. Moreover, similar to that showed in Fig. 9(b), if the charger and QB are in the initial states $|11\rangle$ and $|00\rangle$ (i.e., $n=2$ and $m=2$), respectively, the maximum ergotropy $\bar{\mathcal{E}}$ also corresponds to the second instead of the first dynamical maximum of $\mathcal{E}$ when $R\gtrsim 34$. It is also interesting that the ergotropy $\bar{\mathcal{E}}$ charged on a two-cell QB is much larger than that charged on a single-cell QB when using the same two charging units. This suggests that, in general, the charging process of a QB is extremely complex. It depends on the initial energy and ergotropy in the charger, the residual energy and ergotropy in the QB, as well as the coupling manner and coupling strength between them. To find more intrinsic relations underlying these elements and to identify essential ingredients boosting the charging performance of a QB is an intriguing direction. This will not only help us to have a better understanding on the charging mechanism, but can also help us to develop the high-efficiency QB.
@@
Moreover, it is natural for future work to consider the case that the charger and the QB are coupled to a common reservoir with different strengths, just as the two-qubit case studied in Ref. wireless1 , or further consider the case that there is frequency detuning between the qubits and the cavity field, for which the off-resonant effect on generation of two-qubit entanglement has been studied common1 ; common2 . In addition, one can also generalize the results in this work to the structured reservoir, which acts as a mediator between the charger and QB, with other kinds of spectra (e.g., the sub-Ohmic, Ohmic, and super-Ohmic spectral densities Ohmic ). Of course, in these cases we first need to solve the dynamical equation of the system for a general instead of a special multi-qubit initial state, which may be a very intricate and challenging task.
@@ ACKNOWLEDGMENTS
This work was supported by the National Natural Science Foundation of China (Grant No. 12275212, No. T2121001, No. 92265207, and No. 92365301), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSY008), the Youth Innovation Team of Shaanxi Universities, and Technology Innovation Guidance Special Fund of Shaanxi Province (Grant No. 2024QY-SZX-17).

### v2

@@ IV.3 The case of three charging units
When three qubits in the reservoir are treated as the charger ($n=3$) and the fourth one as a QB, the behaviors of $\mathcal{E}$, $\mathcal{E}_{i}$, and $\mathcal{E}_{c}$ are also structurally similar to those showed in Figs. 2 and 7, so we do not display them here. The charging time $\bar{t}$ and the ergotropy $\bar{\mathcal{E}}$ charged on the QB at $t=\bar{t}$ also show qualitatively a similar $R$ dependence to those showed in Figs. 4 and 8, so we also do not display them here. We only point out here that compared to those achieved with the two charging units, $\lambda\bar{t}$ for both scenarios I and II can be further slightly shortened, while $\bar{\mathcal{E}}$ can also be enhanced to some extent.
Figure 10: Comparison of $\bar{\mathcal{E}}$ charged on the QB with different number $n$ of charging units. The left (right) two panels are plotted with the chargers and QB being in the initial states $|\psi\rangle_{c_{1}\ldots c_{n}}$ and $|0\rangle$ ($|1\rangle^{\otimes n}$ and $|\varphi\rangle_{e_{1}}$), respectively. The first five points in each panel are plotted with $R=0.1$, 1, 2, 3, and 4, respectively.
Before ending this section, we provide some further comparison about the charging of a QB by using different numbers of charging units. First, we present in Fig. 10 a comparison of the ergotropy, where we have denoted by ¯$\bar{\mathcal{E}}_{n=1}$ the ergotropy charged on a QB with $n=1$, and likewise for ¯$\bar{\mathcal{E}}_{n=2}$ and ¯$\bar{\mathcal{E}}_{n=3}$. For scenario I, from Fig. 10(a) one can see that if $c_{1,2}=1$, ¯$\bar{\mathcal{E}}_{n=2}>\bar{\mathcal{E}}_{n=1}$ when $R\lesssim 10.34$, and this region expands with a decrease in $c_{1,2}$. This, together with Fig. 10(c), shows that if the charger energy is not maximal initially, one can enhance the ergotropy of the QB by using multiple charging units. As for scenario II, from Fig. 10(b) one can observe that the region in which ¯$\bar{\mathcal{E}}_{n=2}>\bar{\mathcal{E}}_{n=1}$ shrinks with the increase of $e_{1}$. Although Fig. 10(d) shows that the ergotropy can be slightly enhanced for $n=3$, the enhancement is limited. So in contrast to scenario I, the performance of a QB can be enhanced by using multiple charging units only in the not very strong coupling regime. We have also compared the ergotropy charged on the initially mixed battery state $\rho_{\mathrm{ba}}=\mathrm{diag}\{e_{1},1-e_{1}\}$ with different numbers of charging units, and a qualitatively similar phenomenon to that showed in Fig. 10 is observed.
@@
where $\Delta\bar{\mathcal{E}}$ is the net ergotropy charged on the QB, $\Delta E_{\mathrm{ch}}$ is the energy output from the charger, while $\rho_{\mathrm{ch}}(\bar{t})$ denotes the state of the charger at time $\bar{t}$, and likewise for $\rho_{\mathrm{ch}}(0)$. Based on this definition we performed numerical calculations and the result is shown in Figs. 11. Clearly, $P_{\bar{\mathcal{E}}}$ can be noticeably enhanced by increasing the coupling strength. In particular, for scenario I with a single charging unit, if the QB is in its ground state initially (cf. the circles in the left column of Fig. 11), $P_{\bar{\mathcal{E}}}$ approaches 1 for large enough $R$. Moreover, for the case of not very weak coupling, $P_{\bar{\mathcal{E}}}$ decreases when $n$ increases from 1 to 3, whereas this may not always be the case in the weak coupling regime. For scenario II, as depicted in Figs. 11(c) and (e), if a single charging unit is considered, $P_{\bar{\mathcal{E}}}$ also approaches 1 when $R$ approaches infinity, but in this case its increase with the increase of $R$ is slower than that for scenario I. When the two and three charging units are considered, although $P_{\bar{\mathcal{E}}}$ also increases with the increase of $R$, the asymptotic value at $R\rightarrow\infty$ is obviously less than 1. This implies that a considerable amount of energy output from the charger is leaked into the reservoir in this situation.
@@ V Summary and discussion
In summary, we have studied the wireless charging of a QB via $n$ ($n\geqslant 1$) charging units, where both the charger and the QB interact with a structured bosonic reservoir (which may be implemented, e.g., by an electromagnetic field inside a lossy cavity) and there is no direct coupling between them. Inspired by the fact that in a realistic situation, the charger energy may be not maximal and one rarely waits until a battery runs out before charging, we considered two scenarios of wireless charging, that is, scenario I in which the charger is in a superposition of the excited and ground states and the QB is empty (i.e., fully discharged) initially, and scenario II in which the charger is in its excited state and the QB has residual ergotropy (i.e., partially discharged) initially. For both scenarios, we considered the ergotropy, which quantifies the maximum amount of work a QB could supply during unitary cycles, as a figure of merit for comparing the charging performance. Our results showed that the charging time $\bar{t}$, defined as the interaction time at which the QB reaches its dynamical maximum, decreases with the increase of the coupling strength. The charging time $\bar{t}$ also slightly shortens with an increase in the number of the charging units. The ergotropy $\bar{\mathcal{E}}$ charged on the QB at $t=\bar{t}$, on the other hand, increases with the increasing coupling strength. Specifically, if a single charging unit is used, then in the strong coupling regime, the ergotropy $\bar{\mathcal{E}}$ for scenario I increases with the increasing amount of the initial energy in the charger, while that for scenario II is not very sensitive to the residual ergotropy in the QB. In particular, the QB is almost fully charged when the coupling strength is strong enough. In the weak and moderate coupling regimes, however, the case will be different. The incoherent and coherent contributions to ergotropy are also different in different coupling regimes. Moreover, compared to the case of a single charging unit, the ergotropy $\bar{\mathcal{E}}$ charged by two and three charging units is decreased in the strong coupling regime, whereas in the weak and moderate coupling regimes, it can be enhanced to some extent; in particular, such an enhancement is pronounced for scenario I in which the charger energy is not maximal, and this implies that in this situation the multiple charging units could compensate to some extent the lack of energy in a single charging unit. Finally, we have also considered the charging efficiency which quantifies the proportion of output energy that can be converted into extractable work in the QB. It was found that it also increases with the increase of the coupling strength. If a single charging unit is considered, this efficiency approaches 1 when $R\rightarrow\infty$, while for that of the two and three charging units, it is significantly decreased.
@@
While we considered in this paper the charging of a single-cell QB via $n$ charging units, one may also be concerned with the general case in which an $m$-cell QB is charged via $n$ charging units. Although it is hard to obtain a general result for this case due to the limited computing resource, we performed numerical calculations with $n+m\leqslant 4$ and the results showed that for $m=2$ and 3, the dynamical behavior of $\mathcal{E}$ is structurally similar to that for a single-cell QB, and the maximum ergotropy $\bar{\mathcal{E}}$ charged on a QB can also be noticeably enhanced by increasing the coupling strength, but now it cannot be fully charged, even if $R$ approaches infinity. Moreover, similar to that shown in Fig. 9(b), if the charger and QB are in the initial states $|11\rangle$ and $|00\rangle$ (i.e., $n=2$ and $m=2$), respectively, the maximum ergotropy $\bar{\mathcal{E}}$ also corresponds to the second instead of the first dynamical maximum of $\mathcal{E}$ when $R\gtrsim 34$. It is also interesting that the ergotropy $\bar{\mathcal{E}}$ charged on a two-cell QB is much larger than that charged on a single-cell QB when using the same two charging units.
Figure 12: Scaling behaviors of ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba}})$ and $m\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})$ in terms of the number $m$ of the battery cell, where $R=20$, $c_{1}=0.8$, and the charger and QB are in the initial states $|\psi\rangle_{c_{1}}$ and $|0\rangle^{\otimes m}$, respectively. The four lines from top to bottom in each inset are plotted with $m=1$, 2, 3, and 4, respectively.
For the special case of the initial state $|\psi\rangle_{c_{1}}\otimes|0\rangle^{\otimes m}$, that is, the single charger unit is in the state $|\psi\rangle_{c_{1}}$ of Eq. (11) and the $m$-cell QB is in the state $|0\rangle^{\otimes m}$ initially, one can obtain analytically the ergotropy $\mathcal{E}(\rho_{\mathrm{ba}})$ charged on the $m$-cell QB and the ergotropy $\mathcal{E}(\rho_{\mathrm{ba},l})$ charged on each battery cell (see Appendix A). Based on these analytical results, one can show that in the strong coupling regime, for any fixed $m$, the charging time $\bar{t}$ is the same for different $c_{1}\neq 0$, and it scales as $\ln(\lambda\bar{t})\sim-0.5\ln(m+1)+\ln(\sqrt{2}\pi/R)$. Moreover, as is shown in Fig. 12(a), except for the cases of $m=1$ and 2, the maximal ergotropy ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba}})$ charged on the $m$-cell QB is not a monotonic function of $c_{1}$. As for $m\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})$, from Fig. 12(b) one can note that it is also not a monotonic function of $c_{1}$ when $m\geqslant 2$. In addition, one always has ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba}})>m\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})$ when $m\geqslant 2$, and such a collective effect is rooted in the fact that the product of $m$ passive states may not always be passive entangleu . Finally, as can be seen from Fig. 12, for any $c_{1}\neq 0$ ($c_{1}\neq 0$ and 1), ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba}})$ [$m\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})$] decreases with increasing $m$, and in the large $m$ region, $\ln\bar{\mathcal{E}}(\rho_{\mathrm{ba}})$ and $\ln[m\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})]$ are found to be scaled with different exponents. As a result, if the number of battery cells is very large, the QB cannot be charged by a single charging unit. In the weak coupling regime, an analysis similar to that for $R=20$ shows that the ergotropy also shows a similar scaling behavior, so we do not explicitly present the plots here for the purpose of brevity.
The above results suggest that, in general, the charging process of a QB is extremely complex. It depends on the initial states of the charger and the QB, as well as the coupling manner and coupling strength between them. To find more intrinsic relations underlying these elements and to identify essential ingredients boosting the charging performance of a QB is an intriguing direction. This will not only help us to have a better understanding of the charging mechanism, but can also help us to develop the high-efficiency QB.
@@
Moreover, it is natural for future work to consider the case where the charger and the QB are coupled to a common reservoir with different strengths, just as the two-qubit case studied in Ref. wireless1 , or further consider the case where there is frequency detuning between the qubits and the cavity field, for which the off-resonant effect on generation of two-qubit entanglement has been studied common1 ; common2 . In addition, one can also generalize the results in this work to the structured reservoir, which acts as a mediator between the charger and QB, with other kinds of spectra (e.g., the sub-Ohmic, Ohmic, and super-Ohmic spectral densities Ohmic ). Of course, in all these cases we first need to solve the dynamical equation of the system for the general instead of a special multiqubit initial state, which may be a very intricate and challenging task. Beyond that, it would also be worthwhile to further consider the finite temperature reservoir and explore how it affects the wireless charging of a QB. Intuitively, the increasing temperature may degrade the charging performance as the thermal fluctuations will induce more energy dissipation. But a general study on the details of the finite temperature effects is still needed. Finally, while we take a multimode cavity field as the mediator for energy exchange, another direction for future study may be the Tavis-Cummings model, where the atoms play the role of both charger and battery, and a single-mode cavity field plays the mediated role in the charging process. Similar studies on the Tavis-Cummings QB can be found in Refs. qcqb1 ; shihl3 ; TCqb1 ; noiseqb8 , where the cavity field in various initial states (Fock state, coherent state, squeezed vacuum state, etc.) serves as the charger and $N$ noninteracting two-level atoms serve as the QB.
@@ ACKNOWLEDGMENTS
This work was supported by the National Natural Science Foundation of China (Grants No. 12275212, No. T2121001, No. 92265207, and No. 92365301), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSY008), the Youth Innovation Team of Shaanxi Universities (Grant No. 24JP177), and Technology Innovation Guidance Special Fund of Shaanxi Province (Grant No. 2024QY-SZX-17).
DATA AVAILABILITY
The data that support the findings of this article are openly available data .
Appendix A Solution of the model for a special initial state
When the charger is in the initial state $|\psi\rangle_{c_{1}}$ of Eq. (11) and the $m$-cell QB is in the initial state $|0\rangle^{\otimes m}$, as $|0\rangle^{\otimes(m+1)}$ does not decay in time, the state of the total system (the charger-battery system plus the reservoir) at time $t$ can be written as
$|\Psi(t)\rangle=$ $\sqrt{c_{1}}\Bigg{[}\nu_{1}(t)|1\rangle_{S}|\bar{0}\rangle_{R}+\nu_{2}(t)\sum_{j=2}^{m+1}|j\rangle_{S}|\bar{0}\rangle_{R}$ (A1)
$+\sum_{k}\nu_{k}(t)|0\rangle_{S}|\bar{1}_{k}\rangle_{R}\Bigg{]}+\sqrt{1-c_{1}}|0\rangle_{S}|\bar{0}\rangle_{R},$
where $|0\rangle_{S}=|0\rangle^{\otimes(m+1)}$ denotes the ground state of the charger-battery system, $|\bar{0}\rangle_{R}$ is the vacuum state of the reservoir, $|j\rangle_{S}$ is the state of the charger-battery system with only qubit $j$ being in the excited state, and $|\bar{1}_{k}\rangle_{R}$ is the state of the reservoir with only one excitation in mode $k$. Using the same method as in Ref. common1 , we obtain the coefficients $\nu_{1}(t)$ and $\nu_{2}(t)$ as
$\nu_{1}(t)=\frac{p(t)+m}{m+1},~{}\nu_{2}(t)=\frac{p(t)-1}{m+1},$ (A2)
where the time-dependent parameter $p(t)$ is given by
$p(t)=e^{-\frac{1}{2}\lambda t}\left(\cosh\frac{\zeta t}{2}+\frac{\lambda}{\zeta}\sinh\frac{\zeta t}{2}\right),$ (A3)
with $\zeta=\lambda\sqrt{1-2(m+1)R^{2}}$.
From Eq. (A1), one can obtain the state of the $m$-cell QB as
$\rho_{\mathrm{ba}}=|\xi\rangle\langle\xi|+c_{1}\big{(}1-m|\nu_{2}|^{2}\big{)}|0\rangle_{\mathrm{ba}}\langle 0|,$ (A4)
where
$|\xi\rangle=\sqrt{c_{1}}\nu_{2}\sum_{l=1}^{m}|l\rangle_{\mathrm{ba}}+\sqrt{1-c_{1}}|0\rangle_{\mathrm{ba}},$ (A5)
with $|0\rangle_{\mathrm{ba}}$ and $|l\rangle_{\mathrm{ba}}$ being analogous to $|0\rangle_{S}$ and $|j\rangle_{S}$, respectively, and the only difference is that the former one is for the charger-battery system and the latter one is for the QB. Then, the ergotropy in the $m$-cell QB can be obtained as
$\mathcal{E}(\rho_{\mathrm{ba}})=$ $\omega_{0}\Bigg{[}mc_{1}|\nu_{2}|^{2}$ (A6)
$+\frac{1}{2}\sqrt{1+4mc_{1}^{2}|\nu_{2}|^{2}(m|\nu_{2}|^{2}-1)}-\frac{1}{2}\Bigg{]}.$
Similarly, the state $\rho_{\mathrm{ba},l}$ for the battery cell $l$ ($l=1,\dots,m$) can be obtained as
$\rho_{\mathrm{ba},l}=\begin{pmatrix}c_{1}|\nu_{2}|^{2}&\sqrt{c_{1}(1-c_{1})}\nu_{2}\vspace{0.5em}\\ \sqrt{c_{1}(1-c_{1})}\nu^{*}_{2}&1-c_{1}|\nu_{2}|^{2}\end{pmatrix},$ (A7)
which is the same for different $l$. Hence, the ergotropy in each battery cell can be obtained as
$\mathcal{E}(\rho_{\mathrm{ba},l})=\omega_{0}\Bigg{[}c_{1}|\nu_{2}|^{2}+\frac{1}{2}\sqrt{1+4c_{1}^{2}|\nu_{2}|^{2}(|\nu_{2}|^{2}-1)}-\frac{1}{2}\Bigg{]}.$ (A8)
From Eqs. (A6) and (A8), one can show that if $12(m+1)R^{2}>$ 1, then for any $c_{1}\neq 0$, $\partial\mathcal{E}(\rho)/\partial|\nu_{2}|\geqslant 0$ ($\rho\in\{\rho_{\mathrm{ba}},\rho_{\mathrm{ba},l}\}$), that is, $\mathcal{E}(\rho)$ is always a nondecreasing function of $|\nu_{2}|$. Thus, it takes its maximum when $|\nu_{2}|$ takes its maximum. Based on such an observation, one can obtain the charging time $\bar{t}$ as
$\bar{t}=2\pi/|\zeta|,$ (A9)
which is the same for both $\mathcal{E}(\rho_{\mathrm{ba}})$ and $\mathcal{E}(\rho_{\mathrm{ba},l})$. In the strong coupling region, $\bar{t}$ can be approximated as $\bar{t}\simeq\pi/[(m+1)^{1/2}\Omega]$. By substituting $\bar{t}$ into Eqs. (A6) and (A8), we obtain the maximal ergotropies ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba}})$ charged on the $m$-cell QB and ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})$ charged on each battery cell, respectively.
If $12(m+1)R^{2}<$ 1, both $\mathcal{E}(\rho_{\mathrm{ba}})$ and $\mathcal{E}(\rho_{\mathrm{ba},l})$ take their maxima at $t\rightarrow\infty$, hence the charging time is $\bar{t}\rightarrow\infty$. This corresponds to $p(\bar{t})=0$ and $\nu_{2}(\bar{t})=-1/(m+1)$, and substituting these into Eqs. (A6) and (A8), we obtain the corresponding ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba}})$ and ¯$\bar{\mathcal{E}}(\rho_{\mathrm{ba},l})$, respectively.